September-October Nightly Log files

18 - 19 October 2008 Day 20 - Telescope Commissioning with IRTC
17 - 18 October 2008 Day 19 - Telescope Commissioning with Lucifer
16 - 17 October 2008 Day 18 - Telescope Commissioning with Lucifer
15 - 16 October 2008 Day 17 - Telescope Commissioning with Lucifer
14 - 15 October 2008 Day 16 - Telescope Commissioning with Lucifer - (hotspot: 289,252)
13 - 14 October 2008 Day 15 - Telescope Commissioning
12 - 13 October 2008 Day 14 - LUCIFER Commissioning
11 - 12 October 2008 Day 13 - Lost to bad weather
10 - 11 October 2008 Day 12 - Lost to bad weather
09 - 10 October 2008 Day 11 - LUCIFER Commissioning
08 - 09 October 2008 Day 10 - Rotator hardware failure, lost night
07 - 08 October 2008 Day 9 - On-sky TCS SW testing, Pointing Model
06 - 07 October 2008 Day 8 - On-sky TCS SW testing, M1 Collimation
05 - 06 October 2008 Day 7 - On-sky TCS SW testing, AGw recalibration
04 - 05 October 2008 Day 6 - Lost to bad weather
03 - 04 October 2008 Day 5 - Build 12 TCS Software Testing with dome closed
02 - 03 October 2008 Day 4 - Realignment of WFS Camera with dome closed
1st-2nd October was devoted to LBC Observations because WFS camera still broken.
30 September - 01 October 2008 Day 2 - LBT Lucifer Commissioning Prep
29 - 30 September 2008 Day 1 - LBT Lucifer Commissioning Prep

-- DavidThompson - 24 Nov 2008
Topic revision: r2 - 02 May 2009, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback